The electric field vector E and the magnetic field vector H in


free space are given by £ = 50/p cos(10^6 t + /Bz) a V/m and


H =Ho/p cos( 10^6 t+Bz) ap A/m. Determine their respective


phasor forms and the constants Band Ho such that the fields


satisfy MaxwelJ's equations

1
Expert's answer
2021-11-29T09:16:11-0500

"E=\frac{50}{\rho}\cos(10^{6}t+\beta z)a_{\phi} \frac Vm,"

"H=\frac{H_o}{\rho}\cos(10^{6}t+\beta z)a_{\rho}~ \frac Am,"

"H_o=50,~\beta=1."